Franco and Jason share income and losses in a 2:1 ratio after allowing for salaries of $13,500 and $39,000, respectively. If the partnership suffers a $20,400 loss, by how much would Jason's capital account increase?

a.$27,700

b.$39,000

c.$32,200

d.$14,700
Answer #1
Franco Jason Total
Salary 13500 39000 52500
Loss after salary 72900 =20400+52500
Loss allocated to Jason 24300 =72900*1/3
Jason's capital account increase 14700 =39000-24300
Option D $14,700 is correct